OpenStudy (anonymous):

Total Sales was $960.00 Batteries cost $3.00 and Pens cost $5.00. They sold 5 times more batteries then pens. How many batteries did they sell?

OpenStudy (amistre64):

b(3) + p(5) = 960 b = 5p

OpenStudy (amistre64):

5p(3) + p(5) = 960 15p + 5p = 960 20p = 960 p = 960/20 = 96/2 = 48 b = 5p b=5(48) b = 240 ??

OpenStudy (amistre64):

240(3) + 48(5) = 960 720 + 240 = 960 960 = 960 .... got it :)

OpenStudy (anonymous):

got it. great